# GAPO: Robust Advantage Estimation for Real-World Code LLMs

### v2

@@ Introduction
We evaluate nine diverse instruction-tuned large language models (LLMs), encompassing both general-purpose and code-specialized architectures with parameter counts ranging from 3B to 14B, all sourced from Hugging Face. Since no public dataset captures realistic, history-aware code editing in complex software contexts, we collected an internal dataset of 51,844 real-world code-editing tasks across 10 programming languages, with Go (37.71%), Python (22.14%), and Java (21.03%) dominating the distribution. Each example includes a rich prompt (code context, cursor range, edit history, hints) and the corresponding edited code snippet, as detailed in Table 1. Extensive experiments demonstrate the superiority of our GAPO over both GRPO and DAPO. To sum up, our contributions are:
@@
In code editing tasks, the model is given a lengthy input prompt $q$ that includes the coding context, user prompt, current cursor region, historical edits, and other relevant information. The LLM (with parameters $\theta$) is expected to generate a code snippet $\hat{e}(q,\theta)$ that replaces the current cursor region to fulfill the user’s intent. If the model outputs special tokens indicating “no change,” no edit is applied.
@@
In this study, we focus on code edit in real-world software engineering, where each user maintains complex contextual codebases accompanied by multiple editing histories. These histories capture the users’ individual coding and editing preferences. After an extensive search of publicly available resources, we found no existing dataset or benchmark that adequately represents such realistic code editing behaviors. Therefore, we collected practical code editing data from internal users within our company, reflecting their daily software development activities.
Each data point consists of two fields: <prompt> and <edit>. The <prompt> field contains all necessary inputs, including the code context, user-provided hints, a sequence of edit histories, and the cursor range (i.e., the code segment to be modified). The <edit> field includes the corresponding edited code snippet. The detailed components of each field are summarized in Table 1.
@@
<current code>
<cursor range>
<sequence of edit histories>
<hints>
<edit> <edited code snippet>
Table 1: Components of each data point in the training and evaluation datasets.
The statistics of the training dataset [^2] The evaluation dataset exhibits a similar distribution. are illustrated in Figure 2. The dataset comprises a total of 51,844 code-editing tasks spanning 10 distinct programming languages. The three most represented languages are Go (37.71%), Python (22.14%), and Java (21.03%), which together account for 80.88% of all samples. The remaining languages—C++, Kotlin, TypeScript, JavaScript, C, Rust, and Lua—appear less frequently, with Rust and Lua being the least represented (0.43% and 0.12%, respectively). Overall, the dataset exhibits the following characteristics:
@@
Code editing (li2023codeeditor; guo2024codeeditorbench; nam2025prompting) is more challenging than basic code refinement, as it depends not only on the user prompt but also on the coding context, cursor position, and historical edits. To tackle this challenge, Self-Edit (zhang2023self) fine-tunes a fault-aware neural editor in a generate-and-edit manner, improving code quality and accuracy. EDITLORD (li2025editlord), in contrast, avoids direct prompting or fine-tuning; it first uses an LLM to extract transformation rules from training data and then trains a second LLM to apply these rules to code pairs. IterPref (wu2025iterpref) leverages offline preference learning for iterative debugging, enabling context-aware code editing based on user feedback. LEDEX (jiang2024ledex) automatically collects refinement trajectories and enhances LLMs’ self-debugging ability via supervised fine-tuning and reinforcement learning. Finally, RLEF (gehring2024rlef) performs code editing with real-time execution feedback, optimizing refinement through end-to-end reinforcement learning.

### v3

@@ Introduction
We evaluate nine diverse instruction-tuned large language models (LLMs), encompassing both general-purpose and code-specialized architectures with parameter counts ranging from 3B to 14B, all sourced from Hugging Face. Since no public dataset captures realistic, history-aware code editing in complex software contexts, we collected an internal dataset of 51,844 real-world code-editing tasks across 10 programming languages, with Go (37.71%), Python (22.14%), and Java (21.03%) dominating the distribution. Each example includes a rich prompt (code context, edit history, and code edit range) and the corresponding edited code snippet, as detailed in Table 1. Extensive experiments demonstrate the superiority of our GAPO over both GRPO and DAPO. To sum up, our contributions are:
@@
In code editing tasks, the model is given a lengthy input prompt $q$ that includes the coding context, historical edits, current edit region, cursor position, user prompt, and other relevant information. The LLM (with parameters $\theta$) is expected to generate a code snippet $\hat{e}(q,\theta)$ that replaces the current edit region to fulfill the user’s intent. If the model outputs special tokens indicating “no change,” no edit is applied.
@@
In this study, we focus on code edit in real-world software engineering, where each user maintains complex contextual codebases accompanied by multiple editing histories. These histories capture the users’ individual coding and editing preferences. After an extensive search of publicly available resources, we found only zeta [^2] https://huggingface.co/datasets/zed-industries/zeta that captures realistic code-editing behaviors, but it contains only a few hundred examples. Therefore, we collected a large-scale dataset of more than 50,000 practical code-editing instances from internal users within our company, reflecting their real daily software development activities.
Each data consists of two fields: <prompt> and <edit>. The <prompt> field contains all necessary inputs, including the code context, a sequence of edit histories, the code edit range (with the cursor’s position), and user-provided hints. The <edit> field is the ground truth. The detailed components of each field are summarized in Table 1.
@@
<current code>
<sequence of edit histories>
<code edit range> & <cursor>
@@
<edit> <ground truth>
Table 1: Components of each data in the training and evaluation datasets.
The statistics of the training dataset are illustrated in Figure 2. The dataset comprises a total of 51,844 code-editing tasks spanning 10 distinct programming languages. The three most represented languages are Go (37.71%), Python (22.14%), and Java (21.03%), which together account for 80.88% of all samples. The remaining languages—C++, Kotlin, TypeScript, JavaScript, C, Rust, and Lua—appear less frequently, with Rust and Lua being the least represented (0.43% and 0.12%, respectively). Overall, the dataset exhibits the following characteristics:
@@
Code editing (li2023codeeditor; guo2024codeeditorbench; nam2025prompting) is more challenging than basic code refinement, as it depends not only on the user prompt but also on the coding context, historical edits, edit region, and the cursor position. To tackle this challenge, Self-Edit (zhang2023self) fine-tunes a fault-aware neural editor in a generate-and-edit manner, improving code quality and accuracy. EDITLORD (li2025editlord), in contrast, avoids direct prompting or fine-tuning; it first uses an LLM to extract transformation rules from training data and then trains a second LLM to apply these rules to code pairs. IterPref (wu2025iterpref) leverages offline preference learning for iterative debugging, enabling context-aware code editing based on user feedback. LEDEX (jiang2024ledex) automatically collects refinement trajectories and enhances LLMs’ self-debugging ability via supervised fine-tuning and reinforcement learning. Finally, RLEF (gehring2024rlef) performs code editing with real-time execution feedback, optimizing refinement through end-to-end reinforcement learning.